fairy:

What formula can be used to find the length of a segment in a coordinate plane? and how would u say it in a sentence form ? e-e

steve816:

The distance formula

steve816:

\[\Large d=\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\]

Thronior:

for a segment AB AB^2 = Δx^2 + Δy^2 so AB = squareroot(Δx + Δy) Δx(Delta x) is always the difference in two values x1 and x2 so in this case Δx = xA - xB and Δy = yA - yB

wo1f0mon:

The distance formula is derived from the Pythagorean theorem. To find the distance between two points (x1,y1) and (x2,y2),

HuskyNation:

yup ^

joshiscool:

XD